OpenStudy (anonymous):

Find the x-intercept of f(x)=3x^2-21

OpenStudy (anonymous):

make this equation look like: y=3x^2-21. Then to find the x-intercept, you set y=o. So, 0=3x^2-21. 7=x^2. x=plus/minus sqrt(7).

OpenStudy (anonymous):

so the naswer is y= -21?

OpenStudy (anonymous):

set \[3x^2-21=0\] \[3x^2=21\] \[x^2=7\] \[x=\pm\sqrt{7}\]